# Event-Based Simulation of Stochastic Memristive Devices for Neuromorphic Computing

### v1

@@ III-A1 Calculation of the Equilibrium Point (Metastable Point)
Given the forward switching rate, we can calculate the equilibrium number of switches in the low resistance state at a given time by equating $\lambda_{\text{down}}$ to $\lambda_{\text{up}}$. The expression for this is given in Equation 7, below. A point of dynamic equilibrium is attained when the rate $\lambda_{\text{down}}=\lambda_{\text{up}}$; that is, the rate of switching from the high resistance state to the low resistance state is equal to the rate of switching in the opposite direction. In the case that we have a number of states greater than 2, the determination of the equilibrium state is less straightforward and the steady state vector of the associated Markov state transition matrix must be identified.
@@
III-A2 Additional Discussion
Impact of the Paramter $N$
One strength of a discrete modelling strategy, with a finite number of possible system states (in our case, $N$), is that simulations can be conducted at different resolutions, depending on precision requirements, and the scale of interest. For example, if a memristive device is to be used as a binary storage device, where only the states $1$ and 00 are distinct, then a value of $N=2$ suffices to model the readout value of interest. In contrast, for a simulation where multiple stable states are to be for information storage, a larger value of $N$ may be required. The “best” value of $N$ to use is a matter of perspective - it depends on the extent to which regions of device resistance or behaviour are considered to be equivalent for the intended purpose, as exemplified by the binary storage example, where the many possible configurations of the device are summarised by a single bit (2 possible values) of information.
As a consequence of the fact that the variance and mean of a Poisson process are equal to the rate of the process, the magnitude of the parameter $N$ will scale the expected value and the variance of the final state in a linear fashion. As a result, the standard deviation in the final state, after a given delay, is proportional to the square root of $N$. Thus, we can achieve less stochastic switching by scaling up $N$. Achieving the same mean, for a given delay, for the readout variable can be achieved by applying a straightforward linear scaling of the parameter $n$ in the readout equation. When we increase N, the average jump time decreases linearly with N (the rates of up and down jumps increase proportional to N). This means that over the timescale of interest, many more trials occur, with each trial also contributing to a smaller percentage change in the resistance, resulting in a decrease in the variance with increasing $N$. Continuous and deterministic switching behaviour can be approximated in the limit as $N\rightarrow\infty$.
Error Compared to Timestep-Based Method
Note that the error, in contrast to the GMSM, is related to the level of quantisation and the type of quantisation. In the case of the previous method, we have errors related to missing switching events, which impact the instantaneous probability distribution. We also have error associated with the approximation of a sum of binomials by a normal distribution. The normal distribution’s parameters at each point in time will change with each switching event; however, this is not reflected by the choice of the use of a constant delay between each distribution update. The duration of the delay chosen will have an impact on the accuracy as well.
In contrast, in the case of event-based switching, no error is introduced if the parameter N accurately reflects the physical mechanism. Using a value of N that is smaller will introduce a quantisation error (but no bias) to the switching process evaluation. That is - the expected value of the process after given number of switching events will be the same, since the value of $n_{eq}$ scales linearly with $N$, and the associated parameters of the readout equation can be adjusted accordingly.
Note that integrate and fire sampling for continuous functions of time, as described in Section II-F3, will unavoidably introduce discretisation errors. If the functions vary slowly in time, then the number of additional events due to the introduction of additional functions will be small. However, if the functions vary quickly in time and the number of generated events is large, then the benefits of event-based modelling approaches diminish, and the model’s events become dense in time. This still, however, offers benefits over a uniform sampling approach, since it associates the density of samples in time with the rate of information change that impact the model updates (and hence the error associated with discretisation in time).
IV RRAM Metal Oxide Device Example
In this section, we look at an example of a titanium dioxide memristive device to model the resistive drift phenomenon under zero bias.
IV-A Physical Model
Modelling of metal oxide metal-insulator-metal (MIM), and more specifically - Titanium Dioxide MIM, devices has been studied previously. Titanium Dioxide memristor switching has been previously modelled as a single conductive pathway of varying resistance [5], and other similar devices, such as tantalum oxide memristors have been modelled as exhibiting the dominance of a single filament or conductive region [17].
Based on initial investigation of a drift dataset (see Section IV-C below), we model the conductance of the device as linearly proportional to state variable $n$. Based on this information, we hypothesise that the mechanism underlying the switching of the device modulates the width of a conductive region within the active substrate, bridging a connection betwen the electrodes. Phenomena that could account for this behaviour include a single conductive filament that grows or reduces in width, or multiple conductive filaments that each contribute linearly to the conductance of the device. We observe that it is necessary to use a value of $n_{\text{extra}}$ and $m_{\text{extra}}$ much greater than the value of $N$. This suggests a thresholding effect, where the rate of switching is relatively constant in the switching region, under a constant voltage bias. This corresponds somewhat to the ideal behaviour for a memristor, which is why such titanium dioxide devices were presented as candidate memristive devices originally [5]. We assume a large $n_{\text{extra}}=10\times N$ such that the switching rates remain relatively constant across the considered range of resistance values.
IV-B Readout Equation
We can also invert the equation to find $R$ in terms of $n$:
IV-C Dataset
To evaluate the behaviour of the model under controlled conditions, we created a dataset based on the evolution of memristive states. We collected 2751 device hours of data across two separate memristor arrays. We set each of the devices in the array to a random initial resistance and measured the evolution of the resistance over time, using a short 30ms pulse of 0.3V magnitude to read the instantaneous resistance every 10 seconds. In total, 208 data series were collected, of varying length. The devices used were thin-film Titanium Dioxide ($TiO_{2}$) memristors with $Al_{2}O_{3}$ interstitials (the same as [5]), placed on the diagonal of crossbar arrays to mitigate sneak path currents. The measurements were collected using the ArCONE memristor characterisation platform [18].
After collection, the series were inspected and a data cleaning step was performed. As part of this step, series for memristors which had failed or degenerated to a non-volatile state after programming, as well as clearly anomalous measurements within series due to measurement errors, were removed. Following data cleaning, 14 series were excluded from the dataset, leaving 196 series.
IV-C1 Quantisation
$\mathcal{R}_{\text{bound}}=$$\{-\infty,R(n_{\text{thresh}}+0.5),$
$R(n_{\text{thresh}}+1.5),\cdots,R(N-0.5),\infty\}$
$\mathcal{R}_{\text{val}}=$$\{R(n_{\text{thresh}}),R(n_{\text{thresh}}+1),$
$\cdots,R(N-1),R(N)\}$
IV-C2 Pairs
IV-D Model Fitting
IV-D1 Linear Conductance Approximation
@@
For the chosen timestep of $10,000$ seconds, we find a value of $a=3.62$ yields the best fit, for the given parameters $g_{\text{diff}}$, $g_{\text{parallel}}$, $N$, and $n_{\text{thresh}}$. This associated line of best fit for this linear conductance change model is shown in Figure 7. The moving averages are plotted to enable a clearer view of the general trend. Although their shape is clear, they do include bias associated with the (symmetric) averaging window size, with a larger window size tending to overestimate the mean change, as can be seen from the larger magntiude of the moving average plot for a window size of 50, compared to the model fit to the data. The moving average window of size 5 yields a (stochastic) function whose mean closely matches the model.
IV-D2 Parameter Fitting and Constant Rate Approximation
@@
For the given value of the parameter $a$, we can determine an appropriate value of $V_{a}$. We first select a value of $V_{\text{off}}$ to ensure an equilibrium point (see Section III-A1) that is far above the active switching region (in terms of resistance), to ensure that the assumption that the rate is constant over the considered range holds, since the equilibrium point at a particular bias sets the zero-crossing point for the rate equation, hence our constant rate assumption will break down as the value of $n$ approaches $n_{\text{eq}}$. In our case, this corresponds to setting an equilibrium point with a value of $n<<n_{\text{thresh}}$. We can see in Figure 8 how the choice of $V_{\text{off}}$ impacts $n_{\text{eq}}$; and hence, the constant rate assumption, and the fit to the linear conductance model. Assuming a constant rate and a given value of $V_{\text{off}}$, we can derive a value of $V_{a}$ according to the following equation:
@@
IV-D3 Volatility
In the case of thin film memristors, volatility can be decomposed into two variables. The first of these is the temperature (that may increase due to Joule heating) and the second is a dimensionless volatility state variable to account for additional volatile effects, such as disruptions to the internal structure of the memristor on application of a bias voltage and subsequent restoration of order following its removal (interfacial entropy minimisation).
IV-E Simulation
In addition to demonstrating the fit of the zero-bias mean values to those determined according to the linear conductance model, we demonstrate the capabilities of the model to simulate memristive behaviour through a series of simulated experiments for a titanium dioxide thin film memristive device.
IV-E1 Experimental Parameters
The parameters used for the experiments, chosen for $\Delta T=10000$, are summarised in Table I. The choices for these parameters are given in Sections IV-C1 and IV-D2.
TABLE I: Table of parameters for the model, for titanium dioxide memristive devices.
@@ IV-E1 Experimental Parameters
IV-E2 Switching Experiments
We use the model, with the derived parameters, to generate example device data for simulated input signals. Our first experiments assume there is no volatile behaviour. In the second set of experiments, we assume Joule heating - modulated by the current flowing through the device - and associated restoration of the equilibrium temperature according to Newton’s law of cooling, as described in Section IV-D3.
We perform three sets of experiments: pulsing with positive voltage pulses (increasing/decreasing in magnitude) to increase the device resistance, and with negative voltage pulses (increasing/decreasing in magnitude) to decrease the device resistance. We also examine the effect of alternating pulses. The results of these experiments are shown in Figure 10.
The input signal parameters are given in Table II. We simulate series from starting resistances ranging between $10k\Omega$ and $100k\Omega$. We choose the parameters for the input signals in order to highlight model behaviour. In all the experiments, we see that the model is biased towards movement to a higher resistance state, with negative pulses having a lesser impact on reducing the resistive state than positive pulses have on increasing the state. During the inactive period of each pulse, we see the device’s resistance drift in the upwards direction as a result of the impact of $V_{\text{off}}$.
TABLE II: Parameters for the input signals in the simulated switching experiments. In the case of the positive descending experiment, we observe that too large a signal amplitude drives the memristor immediately to a high resistance state, so we lower the initial amplitude to better observe the behaviour. In the case of the alternating experiment, we use a small duty cycle such that the active pulse duration is 1s.
@@ IV-E2 Switching Experiments
Figure 10: Simulated experiments for a variety of different stimulating voltage inputs. From the alternating pulses experiment, it is clear that the preferred direction of movement (the equilibrium resistance) being a high value, results in the final resistance being higher than the initial resistance, due to the impact of resistive drift, despite the same number and magnitude of positive and negative pulses being applied. Green lines indicate voltage inputs and red/blue lines are resistance evolutions.
V Conclusions and future work

### v2

@@ III-F1 Calculation of the Equilibrium Point (Metastable Point)
The change of the memristive state in the absence of a voltage bias is usually referred to as the phenomenon of resistive drift, or volatility. Given the forward switching rate, we can calculate the equilibrium number of switches in the low resistance state at a given time by equating $\lambda_{\text{down}}$ to $\lambda_{\text{up}}$. The expression for this is given in Equation 10, below. A point of dynamic equilibrium is attained when the rate $\lambda_{\text{down}}=\lambda_{\text{up}}$; that is, the rate of switching from the high resistance state to the low resistance state is equal to the rate of switching in the opposite direction. In the case that we have a number of states greater than 2, the determination of the equilibrium state is less straightforward and the steady state vector of the associated Markov state transition matrix must be identified.
@@
III-F2 Linear Conductance Approximation
@@
For the chosen timestep of $10,000$ seconds, we find a value of $a=3.62$ yields the best fit, for the given parameters $g_{\text{diff}}$, $g_{\text{parallel}}$, $N$, and $n_{\text{thresh}}$. This associated line of best fit for this linear conductance change model is shown in Figure 6. The moving averages are plotted to enable a clearer view of the general trend. Although their shape is clear, they do include bias associated with the (symmetric) averaging window size, with a larger window size tending to overestimate the mean change, as can be seen from the larger magnitude of the moving average plot for a window size of 50, compared to the model fit to the data. The moving average window of size 5 yields a line whose mean closely matches the model.
III-F3 Parameter Fitting and Constant Rate Approximation
@@
For the given value of the parameter $a$, we can determine an appropriate value of $V_{a}$. We first select a value of $V_{\text{off}}$ to ensure an equilibrium point (see Section III-F1) that is far above the active switching region (in terms of resistance), to ensure that the assumption that the rate is constant over the considered range holds, since the equilibrium point at a particular bias sets the zero-crossing point for the rate equation, hence our constant rate assumption will break down as the value of $n$ approaches $n_{\text{eq}}$. In our case, this corresponds to setting an equilibrium point with a value of $n<<n_{\text{thresh}}$. We can see in Figure 7 how the choice of $V_{\text{off}}$ impacts $n_{\text{eq}}$; and hence, the constant rate assumption, and the fit to the linear conductance model. Assuming a constant rate and a given value of $V_{\text{off}}$, we can derive a value of $V_{a}$ according to the following equation:
@@
III-G Simulation
In addition to demonstrating the fit of the zero-bias mean values to those determined according to the linear conductance model, we demonstrate the capabilities of the model to simulate memristive behaviour through a series of simulated experiments for a titanium dioxide thin film memristive device. Due to the event-driven nature of the state transition model we conduct our simulations in this section and in Section IV in Python, rather than a circuit simulation language such as SPICE.
III-G1 Experimental Parameters
The parameters used for the experiments, chosen for $\Delta T=10000$, are summarised in Table II. The choices for these parameters are given in Sections III-E1 and III-F3.
@@ III-G1 Experimental Parameters
TABLE II: Table of parameters for the model, for titanium dioxide memristive devices.
III-G2 Switching Experiments
We use the model, with the derived parameters, to generate example device data for simulated input signals. Our first experiments assume there is no volatile behaviour. In the second set of experiments, we assume Joule heating - modulated by the current flowing through the device - and associated restoration of the equilibrium temperature according to Newton’s law of cooling, as described in Section III-D.
[…]
The input signal parameters are given in Table III. We simulate series from starting resistances ranging between $10k\Omega$ and $100k\Omega$. We choose the parameters for the input signals in order to highlight model behaviour. In all the experiments, we see that the model is biased towards movement to a higher resistance state, with negative pulses having a lesser impact on reducing the resistive state than positive pulses have on increasing the state. During the inactive period of each pulse, we see the device’s resistance drift in the upwards direction as a result of the impact of $V_{\text{off}}$.
@@ III-G2 Switching Experiments
TABLE III: Parameters for the input signals in the simulated switching experiments. In the case of the positive descending experiment, we observe that too large a signal amplitude drives the memristor immediately to a high resistance state, so we lower the initial amplitude to better observe the behaviour. In the case of the alternating experiment, we use a small duty cycle such that the active pulse duration is 1s.
Refer to caption Refer to caption
[…]
@@
Figure 10: Simulated experiments for a variety of different voltage inputs. We see that the positive/negative pulsing experiments tend to bring about switching in the high/low resistance directions, respectively. In the alternating pulses experiment, the high equilibrium resistance results in the final resistance being higher than the initial resistance, due to resistive drift. Green lines are voltage inputs and red-blue lines are resistance series, with bluer lines indicating a lower initial resistance.
IV Neuromorphic Applications
There are a number of ways that the volatility state variable in the modelling framework developed may be used to shed light on approaches that aim to replicate the behaviour of biological neurons using memristors in neuromorphic computing settings. Here, we use the titanium dioxide model realisation we developed in the previous section to elucidate this.
IV-A Stable and Unstable Memristors
We may consider two types of memristor: those which are highly unstable and those that are stable. For the model as presented in Section III, the stability of the memristor can be characterised by the rate of attainment of its equilibrium point, which is determined in large part by the magnitude of the energy barrier $E_{a}$. Unstable devices may be useful for the emulation of neuronal body dynamics, where the interplay between volatility and the application of a stimulating input (for example, current spikes) can result in self-limiting spiking responses. In this context, resistive drift happening over short timescales is analogous to the leakage of membrane potential. In contrast, stable memristors may be more useful for the emulation of synaptic weights. In this context, we can model a combination of short and long term memory effects: Short Term Potentiation (STP) can arise for the programming of a stable memristor to a particular state while the volatility is high. The volatility may cause the state variable to partially revert to its equilibrium point, but eventually dissipates, resulting in stability of the resistive state ( Long Term Potentiation (LTP)). For stable memristors, resistive drift is analogous to the phenomenon of synaptic weight decay.
IV-B Volatility and its Impact
The inclusion of volatility in a memristive model adds a key characteristic that could prove crucial to the emulation of neuronal dynamics such as metastable leakage terms and spiking behaviour (short-term memory), in addition to more stable and persistent changes to the device state (long-term memory). The volatility has an impact on both the equilibrium point and the rate of attainment of the equilibrium point (the time constant of decay). The impact of the volatility state variable on the equilibrium point for a titanium dioxide model was shown in Section III-F1. For stable memristors, the interplay between changes to both rates of switching in the same direction, caused by volatility, and (biased) changes in opposite direction caused by the application of voltage pulses of a given magnitude and time, can lead to memory effects, where states exhibit a dependence on the relative timing of the pulses. A sequence of identical voltage pulses can result in a very different final metastable resistance depending on their spacing in time. For example, in the context of Joule heating, if the voltage pulses are short enough for adiabatic cooling (or relaxation) to occur, the volatility may not increase sufficiently to result in significant switching behaviour. If, however, voltage pulses are more closely spaced, this can result in a larger attained volatility during their application, resulting in switching. Thus, changes in the memristive state are not only dependent on the absolute energy supplied, but also on the way in which this energy is supplied through the pulse shapes and durations.
IV-C Neuromorphic Simulations
We here perform simulations demonstrating the impact of the volatility state variable on two classes of memristor. We first demonstrate the possibility of frequency-dependent switching for stable memristors, and then demonstrate the generation of action potentials for unstable memristors.
IV-C1 Frequency-Dependent Potentiation
In biological neurons, the generation of action potentials can be dependent on the magnitude of the input, according to the leaky integrate and fire neuron model [20, 21, 22], where only an input of a sufficiently high frequency may evoke an action potential - a spike. As a result of the frequency-dependence of postsynaptic neuronal action potential generation and the mechanism of spike timing dependent plasticity (Hebb’s rule) [23], the potentiation of a neuronal synapse may also, consequently, be frequency-dependent. Frequency coding is one of a number of coding schemes used for the communication and processing of information in the brain [24].
Frequency-dependent potentiation has been demonstrated in diffusive memristive devices. Diffusive memristors operate on the basis of the dispersal and regrouping of metal ion clusters under voltage bias [13]. Frequency of application of voltage pulses has been shown to have an effect on the number of pulses required to turn a device ON in protein nanowire catalysed diffusive memristors [25]. This effect was attributed to the resistance drift, due to diffusive regrouping of $Ag+$ ions in the Electrochemical Metallization (ECM) cell, back to the high resistance state during periods where there was no applied bias, meaning that higher frequency voltage pulses required fewer attempts on average to switch devices over a threshold. The number of pulses (of equal magnitude and duration) required to switch the devices from high resistive states to low resistive states was significantly reduced when the frequency of application was increased from 50Hz to 900Hz. A close to linear trend in the number of pulses required as a function of frequency was observed for frequencies higher than 200Hz, with increasing frequency reducing the number of total pulses required to initiate switching. However, a non-linear trend was observed for sub-100Hz frequencies, where the number of pulses increased rapidly. This could be modelled by a volatility state variable whose time constant of decay was on the order of $1/100Hz=0.01s$. In the context of our event-based model, a device with a moderately stable switching characteristic (whose switching rate declines rapidly following the removal of applied bias) can exhibit such frequency dependent behaviour. This is dictated by the state the device is in, as well as the values $V_{a}$ and $V_{\text{off}}$. In Figure 12, we show how changes to the frequency of programming pulses can result in different degrees of potentiation in memristors due to the aforementioned interplay between the volatility and switching. For these experiments, we use a simulation duration of $T=100s$. We update the volatility state variable ($\rho(t)$), at intervals of 0.1s, with updates being generated according to the scheme described in Section II-D3. In these experiments, we supply the same amount of energy to the devices (the same magnitude of voltage over the same amount of time in total), but demonstrate that if the impact of the volatility state variable is high enough, subsequent state transitions can differ greatly, dependent on the frequency of programming, due to the non-linear acceleration in switching effected by an increase in the volatility state variable when the pulses are closely spaced, as compared to sparse pulses in time. We apply square waves of period $T_{\text{pulse}}=1/f$, where $f$ is the frequency of the applied signal. Each square wave period contains a pulse of length 0.1s, and we supply 5 periods as an input signal in each experiment. As such, the same voltage magnitude is applied for the same total amount of time in each case. We modify the parameter $c_{\text{volatile}}=500$, keeping all other parameters as in Table II.
Figure 11: The impact of $\rho(t)$ on $n_{eq}$, for different bias values. Parameters for the calculation are drawn from Table II. As $\rho(t)$ increases, the equilibrium point becomes closer to $N/2$.
[…]
IV-C2 Spiking Behaviour
For unstable memristors, we explore a possible phenomenon enabled by the presence of volatility in a memristive system: the emulation of action potentials.
A “spike” can be defined in the context of a neuromorphic computing architecture as a short-lived increase (or decrease) in an electronic parameter that can be used to process information, such as the voltage, the current, or the conductance. A spike can be thought of as a self-limiting process: once a spike is initiated, it will effect its increase or decrease in the parameter, an associated effect of which is to cause the increase or decrease to cease after a short time. We can envision the creation of spiking behaviour in real devices, if we are able to engineer such a self-limiting process.
We here give a demonstration of how the principle of volatility can be used in the context of the proposed model to generate such a spike. Let us assume that the circuit consists of a memristor in series with a resistor, with an input signal applied across both the memristor and the resistor. This forms a voltage divider circuit, where the proportion of the applied voltage dropped across the memristor will be proportional to its effective instantaneous resistance. If the device is at a high resistance, the proportion of voltage dropped across the memristor will be large. Assuming a volatility whose equilibrium value is proportional to the voltage dropped across the memristor (for example, one can think of Joule heating), the volatility state variable may increase as a result of the magnitude of the applied bias. An increase in the volatility state variable results in a drop in the equilibrium value of the memristor and an increase in the rate of attainment of this value (the net rate). As a result, the device resistance may drop to match the equilibrium point. However, as the resistance decreases, the voltage drops across the memristor, resulting in a decrease in the volatility, which in turn restores the memristor’s equilibrium point to its initial high resistance value, causing the device to return to the high resistance state. One might potentially use the current output of the circuit in order to read and pass on this spike to other parts of the neuromorphic circuit.
As we can see from Equation 10, the equilibrium state is a function of the volatility. Regardless of the parameters, an increase in the volatility results in a shift in the equilibrium state towards the point $n_{eq}=N/2$, as $\lambda_{\text{up}}$ becomes equal to $\lambda_{\text{down}}$. Since the magnitude of both the up and down switching rates also goes up, the rate of attainment of the equilibrium point also goes up. This is visualised in Figure 11.
Biological neurons exhibit two responses to constant input stimuli: firing latencies or rates that are proportional to the magnitude of the input signal (time and frequency encoding, respectively), and a firing response that diminishes over time if the input signal remains constant (adaptation). We propose a method for producing such behaviour using the event-based model, through the dynamics of the volatility state variable.
We can imagine a memristive device with a natural equilibrium state which corresponds to a low resistance. At the same time, we can imagine a volatility state variable which increases in response to the power dissipated by the device. At low resistances, the power dissipated in the memristor will be high, while at high resistances, the reverse will be true, if we assume the voltage drop across the device to be constant.
If the device is at a low resistance, the dissipation of power will result in an increase in the volatility, which results in an increase in the equilibrium resistance (a drop in $n_{eq}$). As the device resistance rises in response, the power dissipated through the device drops. Due to the time constant of decay for volatility, the drop in the equilibrium state will not be instantaneous, and the device resistance will continue to rise. Finally, once the volatility decays sufficiently, the equilibrium resistance will restore to close to its original value, and the device will decrease in resistance, completing a “spike”. The number of such events that happen per unit time will be related to the magnitude of the applied signal. This is shown in Figure 13. In order to produce such a spike train, several parameters were modified, as given in Table IV. The behaviour is essentially brought about through a controlled oscillatory (self-inhibiting) - or transient - response.
Parameter Value
$g_{\text{step}}$ $10^{-8}$
$g_{\text{parallel}}$ $10^{-6}$
$N$ 1000000
$n_{\text{thresh}}$ 800000
$V_{a}$ 1.0
$V_{\text{off}}$0.8
TABLE IV: Parameters to produce spiking behaviour as shown in Figure 13. Only those differing from Table II are listed.
Figure 13: Spiking current behaviour produced using model parameters that produce oscillatory behaviour. Spiking behaviour can be produced through appropriate parameter selection. The intensity of the spikes, the number of spikes in the train, and the frequency of the resultant spike train are dependent on the magnitude of the applied input.
This example demonstrates how through consideration of the volatility state variable as a modelling tool, with its impact on the equilibrium state of the memristor and the rate of attainment of that state, we can predict memory-dependent, self-limiting, short-term behaviour such as spiking, on simulated memristors, offering motivation for its implementation using real memristive devices in a neuromorphic circuit, through the tuning of device parameters.
V Conclusion